# Inconsistency thresholds revisited: The effect of the graph associated with incomplete pairwise comparisons

### v1

@@ Abstract
The inconsistency of pairwise comparisons remains difficult to interpret in the absence of acceptability thresholds. The popular 10% cut-off rule proposed by Saaty has recently been applied to incomplete pairwise comparison matrices, which contain some unknown comparisons. This paper revises these inconsistency thresholds: we uncover that they depend not only on the size of the matrix and the number of missing entries, but also on the undirected graph whose edges represent the known pairwise comparisons. Therefore, using our exact thresholds is especially important if the filling in patterns coincide for a large number of matrices, as has been recommended in the literature. The strong association between the new threshold values and the spectral radius of the representing graph is also demonstrated. Our results can be integrated into software to continuously monitor inconsistency during the collection of pairwise comparisons and immediately detect potential errors.
@@ 1.1 The need for inconsistency thresholds in the incomplete case
Pairwise comparison matrices have a natural generalisation by allowing for missing entries (Harker,, 1987). Incomplete pairwise comparison matrices extend the range of potential applications to hundreds, or even millions of alternatives. For example, the pairwise comparisons of contestants in sports are usually intransitive (van Ours,, 2024, 2025) and may be easily incomplete if some pairs of contestants do not play against each other. Such case studies are presented by Bozóki et al., (2016); Chao et al., (2018); Csató, (2013); Petróczy and Csató, (2021); Temesi et al., (2024), among others. Analogously, neither completeness nor consistency holds for bilateral remittances (Petróczy,, 2021) or students’ preferences (Csató and Tóth,, 2020)
@@ 1.3 A motivating example
According to Ágoston and Csató, (2022, Table 2), the random index $\mathit{RI}\approx 0.306$ if there are $n=4$ alternatives and $m=2$ missing elements. However, Table 1 reveal that the random index is $\mathit{RI}\approx 0.265$ if the positions of the two missing entries (they are placed in different columns and rows) are taken into account, too.
@@ 2.1 Inconsistency and incomplete pairwise comparison matrices
Incomplete pairwise comparison matrices allow for some missing comparisons. Therefore, $\mathbf{A}=\left[a_{ij}\right]$ is an incomplete pairwise comparison matrix if $a_{ij}>0$ or $a_{ij}=\star$ such that $a_{ij}>0$ implies $a_{ji}=1/a_{ij}$ and $a_{ij}=\star$ implies $a_{ji}=\star$. The number of missing comparisons above the diagonal is denoted by $m$, hence, $\lvert(i,j):a_{ij}=\star\rvert=2m$. Incomplete pairwise comparisons have a natural graph representation. Let $\mathbf{A}$ be an incomplete pairwise comparison matrix. The associated undirected graph is $G=(V,E)$, where the vertices correspond to the alternatives and $e_{ij}\in E$ is an edge if and only if the pairwise comparison $a_{ij}$ is known.
@@
Understanding the properties of the spectral radius is an active research field in discrete mathematics (Stevanović,, 2018). Denote the maximum and minimum degree of a graph $G$ by $\Delta(G)$ and $\delta(G)$, respectively. Hong, (1993) asked the following question more than three decades ago: If $G$ has the smallest possible spectral radius among all graphs with $n$ vertices and $e$ edges, does $\Delta(G)-\delta(G)\leq 1$ hold, namely, is $G$ almost regular? The answer seems to be positive, at least for dense graphs with $e\geq(n-1)(n-2)/2-2$ (Cioaba et al.,, 2024). Analogously, there are recent results on upper bounds of the spectral radius (Guo et al.,, 2019; Jin et al.,, 2024).
@@ 3 Results
Section 3.1 provides a comprehensive analysis in the case of four alternatives, and Section 3.2 focuses on pairwise comparison matrices with two missing entries. Finally, matrices of size five and six are discussed in Section 3.3.
@@ 3.1 Incomplete matrices of size four
If $m=2$, there are two possible graphs: (1) the two missing edges are independent, they do not share any vertex (Figure 1); (2) the two missing edges have a common vertex. It is easy to check that the probability of the first case (1) equals
$\frac{n(n-1)/2-(n-1+n-2)}{n(n-1)/2-1}=1-\frac{4n-8}{n(n-1)-2},$
which is only $1/5=20\%$ for $n=4$, but converges to $1$ when $n$ increases.
@@ 3.3 Incomplete matrices of size five and six
If $n=5$, all cases are investigated, that is, the number of missing entries varies from 1 to 5 because $m=6$ implies the existence of a consistent completion (if the unknown entries remain unconstrained). Figure 3 presents the 16 possible graphs if $n=5$ and $2\leq m\leq 5$.
Hence, there are $1+2+4+5+5=17$ random indices, depending on the structure of the graph. All of them are reported in Table A.1 in the Appendix, together with their probability of occurrence if the number of missing entries $m$ is fixed. The probabilities are estimated by simulations, contrary to the exact values derived in Section 3.1. For a fixed value of $m$, the spectral radius is the smallest if the graph is (almost) regular ($G_{2,2}$, $G_{3,3}$, $G_{4,1}$, $G_{5,5}$). These graphs have the smallest probability to occur except for $G_{3,3}$, and the corresponding random index $\mathit{RI}$ is lower than for any other graphs having the same number of edges.
@@ 4 Conclusions
The research is far from closed. First, the random indices are fully computed only up to five alternatives due to the complexity of the numerical calculations. Second, even though the association with the spectral radius is surprisingly strong, other measures of graph “robustness” may provide an even better fit. The association between the random index and the spectral radius may also turn out to be a mere correlation, which is explained by a third, currently hidden indicator.
@@ Acknowledgements
The research was supported by the National Research, Development and Innovation Office under Grant FK 145838 and the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.

### v3

# Refined thresholds for inconsistency: The effect of the graph associated with incomplete pairwise comparisons

@@ Abstract
The inconsistency of pairwise comparisons remains difficult to interpret in the absence of acceptability thresholds. The popular 10% cut-off rule proposed by Saaty has recently been applied to incomplete pairwise comparison matrices, which contain some unknown comparisons. This paper refines these inconsistency thresholds: we uncover that they depend not only on the size of the matrix and the number of missing entries, but also on the undirected graph whose edges represent the known pairwise comparisons. Therefore, using our exact thresholds is especially important if the filling in patterns coincide for a large number of matrices, as has been recommended in the literature. The strong association between the new threshold values and the spectral radius of the representing graph is also demonstrated. Our results can be integrated into software to continuously monitor inconsistency during the collection of pairwise comparisons and immediately detect potential errors.
@@ 1.1 The need for inconsistency thresholds in the incomplete case
Pairwise comparison matrices have a natural generalisation by allowing for missing entries (Harker,, 1987). Incomplete pairwise comparison matrices extend the range of potential applications to hundreds, or even millions of alternatives. For example, the pairwise comparisons of contestants in sports are usually intransitive (van Ours,, 2024, 2025, 2026) and may be easily incomplete if some pairs of contestants do not play against each other. Such case studies are presented by Bozóki et al., (2016); Chao et al., (2018); Csató, (2013); Petróczy and Csató, (2021); Temesi et al., (2024), among others. Analogously, neither completeness nor consistency holds for bilateral remittances (Petróczy,, 2021) or students’ preferences (Csató and Tóth,, 2020)
@@ 1.3 A motivating example
According to Ágoston and Csató, (2022, Table 2), the random index $\mathit{RI}\approx 0.306$ if there are $n=4$ alternatives and $m=2$ missing elements. However, Table 1 reveals that the random index is $\mathit{RI}\approx 0.265$ if the positions of the two missing entries (they are placed in different columns and rows) are taken into account, too.
@@ 2.1 Inconsistency and incomplete pairwise comparison matrices
Incomplete pairwise comparison matrices allow for some missing comparisons. Therefore, $\mathbf{A}=\left[a_{ij}\right]$ is an incomplete pairwise comparison matrix if $a_{ij}>0$ or $a_{ij}=\star$ such that $a_{ij}>0$ implies $a_{ji}=1/a_{ij}$ and $a_{ij}=\star$ implies $a_{ji}=\star$. The number of missing comparisons above the diagonal is denoted by $m$, hence, $\lvert(i,j):a_{ij}=\star\rvert=2m$. Incomplete pairwise comparisons have a natural graph representation. Let $\mathbf{A}$ be an incomplete pairwise comparison matrix. The associated undirected graph is $G=(V,E)$, where the vertices correspond to the alternatives and $e_{ij}\in E$ is an edge if and only if $i\neq j$ and the pairwise comparison $a_{ij}$ is known.
@@
Understanding the properties of the spectral radius is an active research field in discrete mathematics (Stevanović,, 2018). The spectral radius of a $k$-regular graph, where each vertex is connected to the same number of vertices $k$ (has degree $k$), is equal to $k$. Denote the maximum and minimum degree of a graph $G$ by $\Delta(G)$ and $\delta(G)$, respectively. Hong, (1993) asked the following question more than three decades ago: If $G$ has the smallest possible spectral radius among all graphs with $n$ vertices and $e$ edges, does $\Delta(G)-\delta(G)\leq 1$ hold, namely, is $G$ almost regular? The answer seems to be positive, at least for dense graphs with $e\geq(n-1)(n-2)/2-2$ (Cioaba et al.,, 2024). Analogously, there are recent results on upper bounds of the spectral radius (Guo et al.,, 2019; Jin et al.,, 2024).
@@ 3 Results
Section 3.1 provides a comprehensive analysis in the case of four alternatives, and Section 3.2 focuses on pairwise comparison matrices with two missing entries. Matrices of size five and six are discussed in Section 3.3. Finally, Section 3.4 presents an approximation formula based on the number of alternatives, the number of missing elements, and the value of the spectral radius that can be used without costly computations even if $n\geq 7$.
@@ 3.1 Incomplete matrices of size four
If $m=2$, there are two possible graphs: (1) the two missing edges are independent, they do not share any vertex (Figure 1); (2) the two missing edges have a common vertex. The probability of case (1) is
$\frac{n(n-1)/2-(n-2+n-2+1)}{n(n-1)/2-1}=1-\frac{4n-8}{n(n-1)-2},$
since there exist n(n1)/21n(n-1)/2-1 different edges after one is missing, and the missing edge has n2n-2 neighbouring edges at both of its vertices. This value equals only $1/5=20\%$ for $n=4$, but converges to $1$ when $n$ increases.
@@
Considering all possible matrices allows us to exactly compute the difference between Methods 1–3 mentioned in Section 2.2, too. If the missing entries are continuous and not constrained to the interval $\left[1/9;\,9\right]$ (Method 1), a closed formula exists for the optimal completion if $n=4$: Csató et al., (2024) prove that the incomplete eigenvector and the incomplete logarithmic least squares methods (Kwiesielewicz,, 1996; Bozóki et al.,, 2010; Bozóki and Tsyganok,, 2019) imply the same optimal completion, and the optimal completion according to the incomplete logarithmic least squares method can be obtained as a solution of a linear system of equations (Kaiser and Serlin,, 1978; Bozóki et al.,, 2010). Thus, there is no need to use the algorithm suggested by Bozóki et al., (2010, Section 5) to solve optimisation problem (1). The random indices are 0.2528 and 0.2923 for Graphs 1 and 2, respectively, which are reduced by 7.6% and 4.4% compared to the values presented in Table 1.
If the missing entries are chosen from the discrete Saaty scale (Method 3), the random indices are 0.2663 and 0.3174 for Graphs 1 and 2, respectively, which are higher by 0.31% and 0.65% compared to the values presented in Table 1. Therefore, although our continuous relaxation clearly lowers the random index, the extent of this decrease compared to using the discrete scale seems to be minimal.
@@ 3.3 Incomplete matrices of size five and six
If $n=5$, all cases are investigated, that is, the number of missing entries varies from 1 to 5 because $m=6$ implies the existence of a consistent completion (if the unknown entries remain unconstrained). Only one graph exists if $m=1$, when the threshold given in Ágoston and Csató, (2022) remains valid.
@@
Figure 3 presents the 16 possible graphs if $n=5$ and $2\leq m\leq 5$. The $1+2+4+5+5=17$ random indices—depending on the structure of the graph—are reported in Table A.1 in the Appendix, together with their probability of occurrence if the number of missing entries $m$ is fixed. The probabilities are estimated by simulations, contrary to the exact values derived in Section 3.1. For a fixed value of $m$, the spectral radius is the smallest if the graph is (almost) regular ($G_{2,2}$, $G_{3,3}$, $G_{4,1}$, $G_{5,5}$). These graphs have the smallest probability to occur except for $G_{3,3}$, and the corresponding random index $\mathit{RI}$ is lower than for any other graphs having the same number of edges.
@@
Table 3: Spectral radius and the distribution of triads for graphs representing
incomplete pairwise comparison matrices of size $n=6$ with $m=6$ missing entries
Graph Spectral radius Number of triads with missing entries $\mu$
$\mu=0$ $\mu=1$ $\mu=2$ $\mu=3$
$G_{1}$ 3.1819 4 8 8 0
$G_{2}$ 3.2361 4 8 8 0
$G_{3}$ 3.0868 3 10 7 0
$G_{4}$ 3.2814 4 10 4 2
$G_{5}$ 3.1692 3 11 5 1
$G_{6}$ 3.3234 4 10 4 2
$G_{7}$ 3.2227 4 9 6 1
$G_{8}$ 3.1149 3 10 7 0
$G_{9}$ 3.4037 5 8 5 2
$G_{10}$ 3.2361 4 9 6 1
$G_{11}$ 3.3839 5 7 7 1
$G_{12}$ 3.2618 5 6 9 0
$G_{13}$ 3.3539 5 6 9 0
$G_{14}$ 3.2948 4 9 6 1
$G_{15}$ 3.1413 2 14 2 2
$G_{16}$ 3.0922 2 13 4 1
$G_{17}$ 3.1888 3 11 5 1
$G_{18}$ 3 2 12 6 0
$G_{19}$ 3.3723 4 12 0 4
$G_{20}$ 3 0 18 0 2
  • Each incomplete pairwise comparison matrix has 20 triads.
  • The graphs are depicted in Figure 4.
The strong association between the spectral radius of the representing graph and the random index seems to be unexpected. Hence, we provide some theoretical intuition for this relation. Table 3 presents the number of different triads (pairwise comparison matrices of size three) for the 20 graphs with $n=6$ and $m=6$ that can be seen in Figure 4. Six alternatives imply 20 triads in all cases, but these triads are quite different with respect to the number of missing entries $\mu$. For example, in graph $G_{1}$, there are 4 triads without missing entries, and 8-8 triads with one and two missing entries, respectively. On the other hand, graph $G_{20}$ has 18 triads with one missing entry, and 2 triads with three missing entries.
Inconsistency is closely connected to triads since the definition of consistency involves three alternatives. Indeed, several inconsistency indices simply aggregate the inconsistencies of all triads (Brunelli,, 2018). Even though the inconsistency index $\mathit{CI}$ is not such a triad-based index, Cavallo, (2020, Table 2c) uncovers an extremely strong monotonic increasing correlation between $\mathit{CI}$ and the inconsistency index suggested by Peláez and Lamata, (2003), which is the average determinant of all sub-matrices of order three, for $4\leq n\leq 7$. Furthermore, there exists a unique inconsistency index (up to multiplication) if $n=3$ (Csató,, 2019); hence, $\mathit{CI}$ is functionally related to any reasonable inconsistency index in the case of three alternatives.
Thus, reducing $\mathit{CI}$ (the objective function of problem (1)) is essentially equivalent to minimising the inconsistency of the triads. However, the inconsistency of a triad cannot be decreased if it has no missing entries ($\mu=0$), and this is difficult due to the appearance of the same pairwise comparison in n2n-2 different triads if it has one missing entry ($\mu=1$). Consequently, the optimum of optimisation problem (1) is expected to be lower if there are few triads with $\mu=1$ and, especially, $\mu=0$ missing entries.
In order to check the connection between the spectral radius and the distribution of triads, we have estimated two linear regressions by ordinary least squares (OLS). The dependent variable is the spectral radius, while the independent variables are the constant and the number of triads with $\mu=0$ and $\mu=1$ in both cases. The two samples are provided by graphs with $m=6$ (see Table 3) and $m=7$ when the number of alternatives is $n=6$. In both regressions, all coefficients remain significantly positive even at the 0.1% level, and the value of adjusted $R^{2}$ exceeds 0.9. Hence, the spectral radius basically depends on the number of these two types of triads.
To summarise, the value of the spectral radius is probably related to the minimisation of the dominant eigenvalue in (1) through the distribution of triads with respect to the number of missing entries $\mu$. A higher spectral radius is associated with more triads that have at most one missing entry, when decreasing the average inconsistency of triads becomes more challenging. In addition, the optimal value of $\mathit{CI}$ is almost functionally related to the average inconsistency of triads, as shown by Cavallo, (2020).
@@
3.4 An approximation formula for the random index
The computations in Section 3.3 are quite complex and time-consuming, which calls for a heuristic predictor of the random value based on the spectral radius $\rho$.
Let $\mathcal{G}(n,m)$ be the set of graphs for a given number of alternatives $n$ and missing comparisons $m$. Let $\rho_{n,m}$ and $\mathit{RI}_{n,m}$ be the expected values of the spectral radius and random index for a fixed pair of $n$ and $m$, respectively. For any graph $G_{i}\in\mathcal{G}(n,m)$, define
$\Delta\rho_{i}=\rho\left(G_{i}\right)-\rho_{n,m}\text{, and}$
$\Delta\mathit{RI}_{i}=\mathit{RI}\left(G_{i}\right)-\mathit{RI}_{n,m}.$
We estimate the linear regression $\Delta\mathit{RI}_{i}=\beta\Delta\rho_{i}$ by ordinary least squares (OLS) for the 73 observations provided by $n=6$ and $1\leq m\leq 7$ in Table Appendix in the Appendix. This gives $\beta=0.16119$. However, it is not guaranteed that $\beta=0.16119$ remains valid if the number of alternatives is $n\geq 7$. For example, $\beta=0.19219$ if the sample is the set of $5\times 5$ matrices, namely, Table A.1. Indeed, the regression line is apparently steeper in Figure 5.a than in Figure 5.b, but the number of observations is only 17 in the first case. Thus, $\beta=0.16119$ is assumed to hold for any $n\geq 7$ in the following.
Then, for any graph $G_{i}\in\mathcal{G}(n,m)$, the random index $\mathit{RI}_{i}$ is approximated by:
$\mathit{RI}_{i}\approx\mathit{RI}_{n,m}+\beta\Delta\rho_{i}$ (2)
In formula (2), $\Delta\rho_{i}$ depends on $\rho_{i}$ (the spectral radius of graph $G_{i}$) and $\rho_{n,m}$. The latter is estimated by generating 100 thousand random pairwise comparison matrices for each pair of parameters $n$ and $m$, and computing the average spectral radius if the corresponding graph is connected, which is not guaranteed for a high number of missing comparisons $m$. These values are presented in Table Appendix in the Appendix for $6\leq n\leq 9$. Note that this calculation does not require any specific algorithm or solving any optimisation problem.
On the other hand, if $n\geq 7$, the values of $\mathit{RI}_{n,m}$ are neither provided here, nor in Ágoston and Csató, (2022) for all $m$. However, Ágoston and Csató, (2022) present an approximation based on $\mathit{RI}_{n}$ that is available for all plausible number of alternatives $n$ (Aguarón and Moreno-Jiménez,, 2003, Table 1):
$\mathit{RI}_{n,m}\approx\left[1-\frac{2m}{(n-1)(n-2)}\right]\mathit{RI}_{n}.$ (3)
We have checked formula (2) for a pairwise comparison matrix with $n=8$ alternatives, $m=5$ missing entries, and a spectral radius of $\rho_{i}=6.097$. The expected value of the spectral radius is $\rho_{8,5}=5.862$ (see Table Appendix in the Appendix), which leads to $\Delta\rho_{i}=0.235$. The value of $\mathit{RI}_{n,m}$ is estimated from (3) as $1.07$ since $\mathit{RI}_{n}=1.404$. Hence, our heuristic predictor results in $\mathit{RI}_{i}=1.07+0.16119\cdot 0.235=1.1076$. The correct random index, given by the methodology described in Section 2.3, equals $1.117$. Consequently, the error of the approximation is within 1% in this particular case.
@@ 4 Conclusions
The research is far from closed. First, the random indices are fully computed only up to five alternatives due to the complexity of the numerical calculations. Second, even though the association with the spectral radius is surprisingly strong, other measures of graph “robustness” may provide an even better fit.
@@ Acknowledgements
We are grateful to Zsombor Szádoczki for useful advice.
Two reviewers and three anonymous colleagues provided valuable comments and suggestions on earlier drafts.
The research was supported by the National Research, Development and Innovation Office under Grants Advanced 152220 and FK 145838, and the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.
@@
  • If the spectral radius is equal to an integer $k$, the graph is $k$-regular.
@@
Expected value of the spectral radius for graphs representing    incomplete pairwise comparison matrices of size $6\leq n\leq 9$
Alternatives ($n$) Missing comparisons ($m$) Average spectral radius ($\rho_{n,m}$) Connected graphs
[…]
Alternatives ($n$) Missing comparisons ($m$) Average spectral radius ($\rho_{n,m}$) Connected graphs
[…]